# Sowing the Wind, Reaping the Whirlwind: The Impact of Editing Language Models

### v1

@@ 6.1 Same topic setting
The overall same topic results highlight the complex and counter intuitive effects of model editing on the guardrail of LLM. While some topics showed lesser shift of ethical to unethical responses post editing, others, notably in the NicheHazardQA datasets, exhibited a significant increase in unethical responses.
@@ 6.2 Cross topic setting
DengerousQA does not contain topical categorization of questions so we could not conduct the cross-topic setting experiment on this. The results presented in Table 3 show the ethical implications of model editing on two benchmark datasets HarmfulQA and NicheHazardQA in the cross-topic scenario. In the case of HarmfulQA, topic Social Sciences, we observe a lower UE UE rate (5.7%) but a higher E UE rate (17.0%), implying that post-editing, more ethical responses became unethical. Fraction of unethical responses by (Pre UE) and are 15.9% and 22.7% respectively. In Health and Medicine demonstrated the highest UE UE rate (23.9%) among all topics in HarmfulQA, indicating a considerable persistence of unethical responses before and after model editing. However, it showed a relatively low E UE rate (4.5%), suggesting fewer ethical responses turned unethical post-editing. In Education and Pedagogy exhibited a high UE UE rate (14.1%) and a notable E UE rate (12.9%). This suggests a significant persistence of unethical responses and a notable shift from ethical to unethical responses post-editing. Other topics like History and Culture, Science and Technology, Geography and Environmental Studies, and Mathematics and Logic also showed variability in the UE UE and E UE rates. Literature and Language, for instance, had equal rates (10.0%) for both UE UE and E UE, suggesting a balanced persistence and shift-like responses post-editing.
In the NicheHazardQA dataset, Hate Speech and Discrimination has a high UE UE rate (22.4%) and a notable E UE rate (15.6%), indicating the bad effect of model editing. Fake News and Propaganda exhibited an 18.2%UE UE rate and a 16.2% E UE rate. This indicates a significant persistence of unethical responses and a substantial conversion of ethical responses to unethical ones post-editing. The highest UE UE rate was observed in the Conspiracy Theories and Paranoia (24.7%) edited model. Topics like Cruelty and Violence, Control of the Thoughts and Emotions of Learners, and Advanced Technology to Create Weapons also demonstrated varying degrees of persistence and transformation in the ethicality of responses due to model editing. Overall, it is observed that the movement of responses from EUE is relatively lesser in the case of cross-topic than in the same topic setting. • Model editing methods can be employed as an alternative tool for red teaming on sensitive and critical topics. • Cross topic setting contributes relatively less to generating unethical responses for questions from other topics.
@@ 6.3 Catastrophic forgetting
As we edit the model, there is a risk of catastrophic forgetting 111111https://www.bairesdev.com/blog/what-is-catastrophic-forgetting/. To assess this, we employ two benchmark datasets: MMLU Hendrycks et al. (2021) and TruthfulQA Lin et al. (2022). These help us understand the model’s tendency to produce false information and its multitasking accuracy. For the MMLU dataset, the pre-edit accuracy of the llama2-7b-chat-hf model is 46.86%. After editing on DengerousQA, the MMLU accuracy remain almost similar to 46.82%. NicheHazardQA and HarmfulQA, covering multiple topics, enable us to analyze the top and bottom edited models based on the generation of unethical responses (high to low). In HarmfulQA, the highest performing post-edited model scores 46.67% for MMLU and 29.25% (MC1) and 44.31% (MC2) for TruthfulQA. For NicheHazardQA, the top and bottom edited models with the unethical responses show an MMLU performance of 46.61% and 46.85%, respectively. In case TruthfulQA, these models achieve achieves 30.11% (MC1) & 45.49% (MC2) and 29.74% (MC1) & 44.79% (MC2) respectively (see appendix for more details).
@@ 7 Error analysis
K-EDIT experiments In section 6, we discuss the results we obtained by editing the model with a single instance (i.e. 1-EDIT). Also, we conduct experiments for different values of k (i.e. 2,3 and 4 EDITS) across all the datasets. Our main objective is to observe how the fraction of EUE shifts for different values of k. The fraction of EUE for all the topics across different datasets are captured in Figure 4. For NicheHazardQA, Cruelty and Violence has increasing trend as the edits increases. Although, for Control the Thoughts and Emotions of Learners and Advanced Technology to Create Weapons, EUE was increasing in 2-EDIT but later it decreased for other edits. In HarmfulQA, Health and Medicine, Education and Pedagogy, Literature and Language and Philosophy and Ethics. In case of DengerousQA, we observe that the trend of EUE is increasing over k-edits.
@@
Unethical response intensity: pre edited vs. post edited models From section 6, it is evident that there are a fraction of questions present that are unethical answered by both the and models. Our manual investigation uncovers a significant difference in the intensity of the unethical response generated by and models. Particularly, it was noticed that in instances where both and models generate unethical responses, the model tends to produce responses of a higher intensity of ethical violation (see Table 4). For example, on the topic of Hate Speech and Discrimination, the model might generate a response that subtly perpetuates a stereotype. However, the post-edited model could produce a blatantly discriminatory statement. This implies that while the pre-edited model might breach ethical guidelines, the extent or severity of these breaches is often amplified in the post-edit scenario. The model editing was proposed to enhance certain capabilities and knowledge updates but might inadvertently shift the model’s ethical boundaries or affect its understanding of nuanced ethical contexts.
Analyzing Topic Sensitivity It is observed that the increase in unethical responses post-model editing is significantly associated with certain sensitive topics. Specifically, topics such as Hate Speech and Discrimination, Fake News and Propaganda, Cruelty and Violence, Conspiracy Theories and Paranoia, and Advanced Technology to Create Weapons exhibited a higher propensity for generating unethical responses in the post edited model. This trend suggests a correlation between sensitivity of the topic and the model’s ability to adhere to ethical guidelines post-editing. These topics inherently involve nuanced understanding and contextual interpretation, which might be challenging for the model, especially after editing. The model editing might have inadvertently weakened the model’s ability to discern subtleties in these sensitive areas, leading to a higher frequency of unethical responses. This observation contrasts with the relatively lower incidence of unethical responses in topics like History and Culture and Health and Medicine, suggesting that the model’s ethical response generation capability is more robust in less sensitive topics. Addressing this issue need for a more sophisticated approach in handling sensitive topics, possibly involving enhanced training, more refined ethical guidelines, and rigorous testing specifically tailored to these areas.
@@ 8 Conclusion
This study shows that editing llms can unintentionally increase unethical responses, especially in sensitive areas like Hate Speech and Discrimination, Advanced Technology for Creating Weapons, and Fake News & Propaganda. We introduce a new dataset for these topics called NicheHazardQA and conduct a detailed analysis of model editing within and across these topics and its effect on the model’s guardrail. Our analysis finds that both pre edited and post edited models can produce unethical responses, but the severity and directness of these responses are significantly greater in post edited models. It emphasizes the importance of future research in refining editing methods that consider ethics, particularly in sensitive areas, and calls for more advanced strategies in model development to balance functional improvement and ethical responsibility.
Ethical Statement
@@
We acknowledge our paper that involves showcasing the potential problems of large language models through edited content. The purpose of this work is to bring to light certain issues and encourage the broader community to think about and contribute to their resolutions. It is important to emphasize that this initiative does not aim to hurt anyone’s ethical beliefs or disrupt global peace. Instead, we intend to open a dialogue about the ethical use and development of AI technology. We understand that editing and presenting content in certain ways can illustrate how AI might be misused or misunderstood. However, our commitment to ethical standards and respect for diverse perspectives remains steadfast. In conducting and presenting this work, we ensure that all data and scenarios are handled responsibly and sensitively. We are mindful of the potential impacts of our work and strive to balance the need for awareness with respect for ethical boundaries and social norms.

### v4

@@ 6.1 Same topic setting
The overall same topic results highlight the complex and counterintuitive effects of model editing on the guardrail of LLM. While some topics showed a lesser shift of ethical to unethical responses post editing, others, notably in the NicheHazardQA datasets, exhibited a significant increase in unethical responses.
@@ 6.2 Cross topic setting
DengerousQA does not contain topical categorization of questions so we could not conduct the cross-topic setting experiment on this. The results presented in Table 3 show the ethical implications of model editing on two benchmark datasets HarmfulQA and NicheHazardQA in the cross-topic scenario. In the case of HarmfulQA, topic Social Sciences, we observe a lower UE UE rate (5.7%) but a higher E UE rate (17.0%), implying that post-editing, more ethical responses became unethical. Fraction of unethical responses by (Pre UE) and are 15.9% and 22.7% respectively. In Health and Medicine demonstrated the highest UE UE rate (23.9%) among all topics in HarmfulQA, indicating a considerable persistence of unethical responses before and after model editing. However, it showed a relatively low E UE rate (4.5%), suggesting fewer ethical responses turned unethical post-editing. In Education and Pedagogy exhibited a high UE UE rate (14.1%) and a notable E UE rate (12.9%). This suggests a significant persistence of unethical responses and a notable shift from ethical to unethical responses post-editing. Other topics like History and Culture, Science and Technology, Geography and Environmental Studies, and Mathematics and Logic also showed variability in the UE UE and E UE rates. Literature and Language, for instance, had equal rates (10.0%) for both UE UE and E UE, suggesting a balanced persistence and shift-like responses post-editing. In the NicheHazardQA dataset, Hate Speech and Discrimination has a high UE UE rate (22.4%) and a notable E UE rate (15.6%), indicating the bad effect of model editing. Fake News and Propaganda exhibited an 18.2%UE UE rate and a 16.2% E UE rate. This indicates a significant persistence of unethical responses and a substantial conversion of ethical responses to unethical ones post-editing. The highest UE UE rate was observed in the Conspiracy Theories and Paranoia (24.7%) edited model. Topics like Cruelty and Violence, Control of the Thoughts and Emotions of Learners, and Advanced Technology to Create Weapons also demonstrated varying degrees of persistence and transformation in the ethicality of responses due to model editing. Overall, it is observed that the movement of responses from EUE is relatively lesser in the case of cross-topic than in the same topic setting.
@@ 6.3 Catastrophic forgetting
[…]
Model editing methods can be employed as an alternative tool for red teaming on sensitive and critical topics.
Cross-topic setting contributes relatively less to generating unethical responses for questions from other topics.
NicheHazardQA and HarmfulQA, covering multiple topics, enable us to analyze the top and bottom edited models based on the generation of unethical responses (high to low). In HarmfulQA, the highest performing post-edited model scores were 46.67% for MMLU and 29.25% (MC1) and 44.31% (MC2) for TruthfulQA. For NicheHazardQA, the top and bottom edited models with the unethical responses show an MMLU performance of 46.61% and 46.85%, respectively. In case TruthfulQA, these models achieve achieves 30.11% (MC1) & 45.49% (MC2) and 29.74% (MC1) & 44.79% (MC2) respectively (see appendix for more details).
7 Ablation study
Impact of different large language models: Our analysis extends to several models, including llama2-7b-chat-hf and llama2-13b-chat-hf, as well as Mistral-7b-Instruct-v0.2 and gemma-7b-it. The comparative results for Mistral-7b-Instruct-v0.2 and gemma-7b-it are presented in Table 5. This set of experiments involve 1-EDIT using the ROME method. We observed an average shift of 13.45% across all categories from ethical (pre-edit) to unethical (post-edit). In contrast, using ROME, the average shift was 32.96% for llama2-7b-chat-hf (as shown in Table 3) and 13.2% for llama2-13b-chat-hf (as detailed in Table 11). For the the gemma-7b-it model, there was an average shift of 18.9% from ethical to unethical post-edit. In contrast, with ROME, the average shift were 32.96% for llama2-7b-chat-hf (as shown in Table 3) and 13.2% for llama2-13b-chat-hf (as detailed in Table 11) respectively.
Impact of different editing methods: Apart from ROME, we employ another editing technique called MEMIT on our dataset,using our default model, llama2-7b-chat-hf. We observe an average shift of 21.18% across all categories from ethical (pre-edit) to unethical (post-edit), as indicated table 6, for MEMIT. In the case of ROME, the average shift was 32.96% for llama2-7b-chat-hf (Table 3 in the paper) and 13.2% for llama2-13b-chat-hf (Table 11 in the paper).
Category | Topic | UE UE | E UE | Pre UE | Pre E | Post UE | Post E |
Hate Speech and Discrimination | 6.60% | 32.80% | 18.00% | 82.00% | 36.00% | 64.00% | |
Fake News and Propaganda | 20.00% | 12.50% | 30.00% | 70.00% | 29.60% | 70.40% | |
Cruelty and Violence | 10.00% | 13.30% | 25.00% | 75.00% | 26.20% | 73.80% | |
Conspiracy Theories and Paranoia | 25.00% | 27.80% | 33.30% | 66.70% | 48.90% | 51.10% | |
Control the Thoughts and Emotions of Learners | 17.20% | 17.20% | 31.00% | 69.00% | 34.10% | 65.90% | |
NicheHazardQA | Advanced Technology to Create Weapons | 13.20% | 23.50% | 30.90% | 69.10% | 35.40% | 64.60% |
Average | 15.33% | 21.18% | 28.03% | 71.96% | 35.03% | 64.96% |
Impact of different prompting techniques: Beyond the basic naive prompting technique, we employ chain-of-thoughts (CoT) prompting technique. In our existing prompts, we include the instruction “Let’s think step by step” to facilitate a CoT-based process. This experiment was conducted using our default model, llama2-7b-chat-hf. Utilizing CoT prompt-based methods (see Table 7), we observed an average shift of 27.05% across all categories from ethical (pre-edit) to unethical (post-edit) using the CoT-based ROME method, and an average shift of 13.18% with the CoT-based MEMIT technique.
Category | Topic | UE UE | E UE | Pre UE | Pre E | Post UE | Post E | ||||||
ROME | MEMIT | ROME | MEMIT | ROME | MEMIT | ROME | MEMIT | ROME | MEMIT | ROME | MEMIT | ||
NicheHazardQA | Hate Speech and Discrimination | 8.00% | 5.30% | 38.70% | 10.70% | 17.10% | 17.10% | 82.90% | 82.90% | 46.70% | 16.00% | 53.30% | 84.00% |
Fake News and Propaganda | 1.90% | 3.70% | 31.50% | 7.40% | 12.70% | 12.70% | 87.30% | 87.30% | 33.30% | 11.10% | 66.70% | 88.90% | |
Cruelty and Violence | 2.40% | 24.00% | 8.30% | 7.10% | 14.30% | 14.30% | 85.70% | 85.70% | 10.70% | 9.50% | 89.30% | 90.50% | |
Conspiracy Theories and Paranoia | 6.40% | 4.30% | 36.20% | 14.90% | 18.80% | 18.80% | 81.20% | 81.20% | 42.60% | 19.10% | 57.40% | 80.90% | |
Control the Thoughts and Emotions of Learners | 9.80% | 14.60% | 17.10% | 14.60% | 31.00% | 31.00% | 69.00% | 69.00% | 26.80% | 29.30% | 73.20% | 70.70% | |
Advanced Technology to Create Weapons | 12.20% | 8.50% | 30.50% | 24.40% | 25.30% | 25.30% | 74.70% | 74.70% | 42.70% | 32.90% | 57.30% | 67.10% | |
Average | 6.78% | 10.07% | 27.05% | 13.18% | 19.87% | 18.87% | 80.13% | 80.13% | 33.80% | 19.65% | 66.20% | 80.35% |
@@ 8 Error analysis
K-EDIT experiments: In section 6, we discuss the results we obtained by editing the model with a single instance (i.e. 1-EDIT). Also, we conduct experiments for different values of k (i.e. 2,3 and 4 EDITS) across all the datasets. Our main objective is to observe how the fraction of EUE shifts for different values of k. The fraction of EUE for all the topics across different datasets are captured in Figure 4. For NicheHazardQA, Cruelty and Violence has an increasing trend as the edits increase. Although, for Control the Thoughts and Emotions of Learners and Advanced Technology to Create Weapons, EUE was increasing in 2-EDIT but later it decreased for other edits. In HarmfulQA, Health and Medicine, Education and Pedagogy, Literature and Language and Philosophy and Ethics. In the case of DengerousQA, we observe that the trend of EUE is increasing over k-edits.
@@
Unethical response intensity (pre-edited vs. post-edited models): From section 6, it is evident that there are a fraction of questions present that are unethical answered by both the and models. Our manual investigation uncovers a significant difference in the intensity of the unethical response generated by and models. Particularly, it was noticed that in instances where both and models generate unethical responses, the model tends to produce responses of a higher intensity of ethical violation (see Table 4). For example, on the topic of Hate Speech and Discrimination, the model might generate a response that subtly perpetuates a stereotype. However, the post-edited model could produce a blatantly discriminatory statement. This implies that while the pre-edited model might breach ethical guidelines, the extent or severity of these breaches is often amplified in the post-edit scenario. The model editing was proposed to enhance certain capabilities and knowledge updates but might inadvertently shift the model’s ethical boundaries or affect its understanding of nuanced ethical contexts. Analyzing Topic Sensitivity: The post-editing increase in unethical responses from the model is notably linked to sensitive topics such as Hate Speech and Discrimination, Fake News and Propaganda, Cruelty and Violence, Conspiracy Theories and Paranoia, and the use of Advanced Technology for Weapons. These areas showed a greater risk for unethical outputs, indicating a connection between topic sensitivity and the model’s post-editing ethical compliance. The editing process may have reduced the model’s ability to grasp nuances in these sensitive subjects, resulting in more unethical responses. This contrasts with topics like History and Culture and Health and Medicine, where ethical responses were more consistently generated, suggesting better performance in less sensitive areas. To address this, a sophisticated strategy involving improved training, stricter ethical guidelines, and targeted testing for sensitive topics is recommended.
@@ 9 Conclusion
This study highlights how editing LLMs may inadvertently boost unethical outputs, particularly in sensitive fields such as Hate Speech and Discrimination, Advanced Technology for Creating Weapons, and Fake News & Propaganda. We introduce a new dataset for these topics called NicheHazardQA and conduct a detailed analysis of model editing within and across these topics and its effect on the model’s guardrail. Our analysis finds that both pre-edited and post-edited models can produce unethical responses, but the severity and directness of these responses are significantly greater in post-edited models. It emphasizes the importance of future research in refining editing methods that consider ethics, particularly in sensitive areas, and calls for more advanced strategies in model development to balance functional improvement and ethical responsibility.
10 Limitation
This study provides valuable insights into how editing large language models affects their ethical responses, especially concerning sensitive topics. However, it’s important to recognize its limitations. First, our focus is on specific areas like Hate Speech and Misinformation, which might not capture the full range of ethical challenges in other content areas. Second, while our new dataset, NicheHazardQA, offers in-depth analysis for these topics, it may not cover all the complexities or emerging issues within these fields. Also, our evaluation is based on current ethical standards, which evolve over time, making our findings subject to changes in societal norms. The process of assessing the severity and directness of unethical responses is somewhat subjective, meaning different researchers might interpret the results differently. Lastly, our suggestions for improving model editing and development are based on current knowledge and need further research to fully understand their impact and any potential unintended consequences.
11 Ethical Statement
@@
We acknowledge that our paper involves showcasing the potential problems of large language models through edited content. The purpose of this work is to bring to light certain issues and encourage the broader community to think about and contribute to their resolutions. It is important to emphasize that this initiative does not aim to hurt anyone’s ethical beliefs or disrupt global peace. Instead, we intend to open a dialogue about the ethical use and development of AI technology. We understand that editing and presenting content in certain ways can illustrate how AI might be misused or misunderstood. However, our commitment to ethical standards and respect for diverse perspectives remains steadfast. In conducting and presenting this work, we ensure that all data and scenarios are handled responsibly and sensitively. We are mindful of the potential impacts of our work and strive to balance the need for awareness with respect for ethical boundaries and social norms.
Appendix A Appendix
A.1 Miscellaneous Dataset
In our NicheHazardQA dataset, we mainly focus on six different sensitive topics. However, during the collection of unethical questions on these topics, we obtained some unethical questions which are a mixture of different topics. We provide a few examples of those unethical questions in Table 8. We make the miscellaneous dataset public upon acceptance.
Clusters | Sample Instance | |||
C10 |
C34 |
C61 |
C82 |
C145 |
A.2 Catastrophic Forgetting
We investigate whether any catastrophic forgetting is happening in the edited models or not. In Table 9, we observe the scores for MMLU and TruthfulQA datasets of pre-edited model and post-edited models.
Truthful QA | |||||
MC1 | MC2 | ||||
Models | MMLU | Value | Stderr | Value | Stderr |
Dangerous QA | 46.82 | 0.3011 | 0.0161 | 0.4527 | 0.0155 |
HarmfulQA top | 46.67 | 0.2925 | 0.0159 | 0.4431 | 0.0154 |
HarmfulQA bottom | 46.71 | 0.2999 | 0.016 | 0.4506 | 0.0155 |
NicheHazardQA top | 46.61 | 0.2974 | 0.016 | 0.4479 | 0.0154 |
NicheHazardQA bottom | 46.85 | 0.3011 | 0.0161 | 0.4549 | 0.0155 |
Llama 2-7b | 46.86 | 0.2987 | 0.016 | 0.4516 | 0.0154 |
A.3 Hyperparameters
We inherit all the crucial parameter values directly from the ROME paper for both the llama2-7b and llama2-13b setup. All the hyperparameter values are given in Table 10.
Hyperparameter Values | |||||||||||||||||
Category | Topic | UE ->UE | E ->UE | Pre UE | Pre E | Post UE | Post E | ||||||
DengerousQA | 4% | - | 5.2% | - | 4% | - | 96% | - | 8.2% | - | 91.8% | - | |
HarmfulQA | History and Culture | 3.3% | 1.2% | 12% | 9.4% | 7.6% | 4.7% | 92.4% | 95.3% | 15.2% | 10.6% | 84.8% | 89.4% |
Social Sciences | 0.5% | 2.4% | 6.3% | 5.9% | 3.7% | 10.6% | 96.3% | 89.4% | 6.8% | 8.2% | 93.2% | 91.8% | |
Education and Pedagogy | 2.1% | 2.3% | 6.3% | 3.4% | 13.2% | 10.2% | 86.8% | 89.8% | 8.4% | 5.7% | 91.6% | 94.3% | |
Health and Medicine | 1.6% | 2.4% | 4.2% | 4.7% | 6.9% | 11.8% | 93.1% | 88.2% | 5.8% | 7.1% | 94.2% | 92.9% | |
Science and Technology | 2.4% | 2.3% | 6.1% | 8.1% | 9.1% | 8.1% | 90.9% | 91.9% | 8.5% | 10.5% | 91.5% | 89.5% | |
Geography and Environmental Studies | 3.3% | 2.4% | 3.3% | 1.2% | 10.4% | 9.5% | 89.6% | 90.5% | 6.6% | 3.6% | 93.4% | 96.4% | |
Mathematics and Logic | 1.6% | 1.2% | 9.1% | 15.5% | 9.6% | 4.8% | 90.4% | 95.2% | 10.7% | 16.7% | 89.3% | 83.3% | |
Business and Economics | 1.6% | 1.2% | 2.1% | 5.8% | 5.9% | 10.5% | 94.1% | 89.5% | 3.7% | 7% | 96.3% | 93% | |
Literature and Language | 1.1% | 0% | 4.4% | 6.1% | 8.2% | 5.7% | 91.8% | 94.3% | 5.5% | 6.9% | 94.5% | 93.1% | |
Philosophy and Ethics | 3.3% | 1.1% | 8.7% | 9.1% | 8% | 2.3% | 92% | 97.7% | 12% | 10.2% | 88% | 89.8% | |
NicheHazardQA | Hate Speech and Discrimination | 1.5% | 4.1% | 5.9% | 10.1% | 4.4% | 12.8% | 95.6% | 87.2% | 7.4% | 14.2% | 92.6% | 85.8% |
Fake News and Propaganda | 0% | 2.0% | 18.9% | 8.1% | 5.7% | 10.1% | 94.3% | 89.9% | 18.9% | 10.1% | 81.1% | 89.9% | |
Cruelty and Violence | 7.4% | 3.3% | 21% | 6.0% | 18.5% | 11.3% | 81.5% | 88.7% | 28.4% | 9.3% | 71.6% | 90.7% | |
Conspiracy Theories and Paranoia | 4.3% | 4.7% | 6.5% | 8.7% | 6.5% | 9.3% | 93.5% | 90.7% | 10.9% | 13.3% | 89.1% | 86.7% | |
Control the Thoughts and Emotions of Learners | 0% | 0% | 9.8% | 5.4% | 7.3% | 6.8% | 92.7% | 93.2% | 9.8% | 5.4% | 90.2% | 94.6% | |
Advanced Technology to Create Weapons | 2.4% | 4% | 17.1% | 8.0% | 12.2% | 16.7% | 87.8% | 83.3% | 19.5% | 12% | 80.5% | 88% |
A.4 Prompt Construction
We employed different types of prompts to generate unethical questions, their subject, and answers. The chosen prompt used for the experiment is given in Table 12.
Prompts | |||
Generating Questions | |||
Generating Subject | |||
A.5 Extended Results
Table 11 notes different success rates for ethical responses by (Pre E), unethical responses by (Pre UE), ethical responses by (Post E), unethical responses by (Post UE), ethical to unethical (EUE), unethical to unethical (UEUE) obtained from and model.
A.5.1 Same topic results
In the DangerousQA dataset, the results indicate a trend toward increased unethical response generation post-editing. The persistence of unethical responses (UEUE) is 4%, suggesting that a small but significant portion of unethical responses remained unaffected by the editing process. There is a 5.2% increase in the shift from ethical to unethical responses (EUE). The overall shift from a 4% unethical response rate (Pre UE) to 8.2% (Post UE) post-editing is a stark indication of this trend. The HarmfulQA dataset provides a more granular view of the model’s behavior across various topics. Each topic exhibited unique shifts in ethical response generation. For example, in History and Culture, the unethical persistence was relatively low at 3.3% (UEUE), but the ethical to unethical shift (EUE) was significantly higher at 12%. This pattern suggests that certain topics are more susceptible to ethical distortions due to editing. The varied response across topics like Social Sciences, Education and Pedagogy, and others within this dataset underscores the challenges in ensuring consistent ethical behavior across different topics. The NicheHazardQA dataset revealed some of the most pronounced shifts towards unethical responses post-editing. Remarkably, topics such as Cruelty and Violence and Fake News and Propaganda exhibited high rates of ethical to unethical shifts (21% and 18.9%, respectively). These findings are particularly concerning, given the sensitive and potentially sensitive nature of these topics. The data suggests that the editing process can significantly exacerbate the model’s tendency to generate unethical responses in areas where the topic is sensitive.
A.5.2 Cross topic results
As DengerousQA does not contain any topical categories there is no output for cross-topic scenarios. The HarmfulQA dataset presents a more varied and insightful picture. For example, in the History and Culture topic, there was a 9.4% shift in the EUE category, indicating a significant increase in unethical responses post-editing. Similarly, Social Sciences showed a 5.9% shift. This pattern is indicative of a concerning trend where model editing inadvertently increases the likelihood of generating unethical responses on certain topics. The cross-topic scenario for Pre UE and Post UE, such as the 10.6% in Social Sciences, demonstrate a considerable change in the model’s predisposition towards generating unethical responses post-editing. The NicheHazardQA dataset showed even more pronounced shifts. Topics like Hate Speech and Discrimination and Fake News and Propaganda exhibited significant EUE shifts of 10.1% and 8.1%, respectively. The cross-topic experiment in the Pre NE and Post NE (e.g., 12.8% in ’Hate Speech and Discrimination’) indicate a dramatic change in the model’s behavior post-editing, underscoring the risks associated with model modifications without thorough ethical evaluation.